[ In a parallelogram one angle is four times its adjacent angle. Find all the angles of the parallelogram. ]
Answer by TimothyLamb(4379)

---
x = 4y
x + y = 180
---
put the system of linear equations into standard form
---
x - 4y = 0
x + y = 180
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = two opposite angles are 144 degrees each
y = the other two angles are 36 degrees each
---
system matrix
1 -4
1 1
---
inverse of system matrix
0.2 0.8
-0.2 0.2
---
determinant of system matrix = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php